# The Radius Distribution of M dwarf-hosted Planets and its Evolution

### v1

@@ 4.2 Clues to the nature of sub-Neptunes
Our findings, while intriguing, are limited in their statistical robustness not by the precision with which stellar and planet parameters can be determined, but by sample size, the small number of late K and M dwarf systems being a product of the selection criteria for targets of the Kepler prime mission and the telescope’s single fixed pointing. Both the K2 and TESS mission have identified many more transiting planet systems around such stars, some of which have ages assigned by gyrochronology (Gaidos et al., 2023b). However, determining the rotation periods of late-type field stars from K2 or TESS photometry alone remains challenging (Reinhold & Hekker, 2020; Claytor et al., 2022). The PLATO mission and its long-duration (1 year) observation phase fields should yield additional systems with gyrochronologic ages, but only among K dwarfs (Nascimbeni et al., 2022). There are also additional investigations that could be performed with larger samples, for example, the expected dynamical evolution of multi-planet systems due to mutual perturbations (Teixeira & Ballard, 2023).777The orbital period distributions of the young and old sub-samples presented here do not appear significantly different (K-S test ).
| Name | Radius | Age |
| Gyr | ||
| KOI-247.01 | 1.91 +0.10/-0.11 | 1.030.35 |
| KOI-251.01 | 2.65 +0.10/-0.10 | 1.090.20 |
| KOI-251.02 | 0.85 +0.05/-0.05 | 1.090.20 |
| KOI-253.01 | 3.20 +0.13/-0.13 | 6.000.97 |
| KOI-253.02 | 1.85 +0.10/-0.11 | 6.000.97 |
| KOI-254.01 | 11.54 +0.35/-0.35 | 1.140.24 |
| KOI-314.01 | 1.50 +0.05/-0.05 | 2.210.67 |
| KOI-314.02 | 1.33 +0.05/-0.05 | 2.210.67 |
| KOI-314.03 | 0.64 +0.03/-0.03 | 2.210.67 |
| KOI-430.01 | 2.69 +0.10/-0.10 | — |

### v2

@@ 5 Summary and Future Directions
Our findings, while intriguing, are limited in their statistical robustness by sample size rather than the precision with which stellar and planet parameters can be determined. The small number of late K and M dwarf KOIs is the result of the selection criteria for targets of the Kepler prime mission and the telescope’s single fixed pointing. Both the K2 and TESS mission have identified many more transiting planet systems around such stars, some of which have ages assigned by gyrochronology (Gaidos et al., 2023). However, determining the rotation periods of late-type field stars from K2 and especially TESS photometry alone is challenging (Reinhold & Hekker, 2020; Claytor et al., 2022). In the Gaidos et al. (2023) catalog of late K- and early M-type stars with established rotation periods and ages there are only 18 TESS Objects of Interest (TOIs). This situation will improve with the discovery of more TOIs hosted by very cool stars, plus long-period planets near the water-condensation limit detected with only one or two transits by TESS and confirmed by other observatories (Gill et al., 2020; Cañas et al., 2020; Yao et al., 2021; Magliano et al., 2024). However, obtaining rotation periods of middle-aged cool host stars for gyrochronology will in most cases require sources of photometry other than TESS or analysis of activity indicators in the spectra obtained for radial velocity monitoring. The PLATO mission and its long-duration (two-year) fields could yield additional systems with gyrochronologic ages, depending on target selection (Nascimbeni et al., 2022).
The water condensation limit should be marked by a transition in the abundance of H2O in the upper atmospheres of cool sub-Neptunes, a boundary that is beginning to be explored by infrared transmission spectroscopy by JWST, e.g. TOI-270d (Holmberg & Madhusudhan, 2024). There are also additional investigations that could be performed with larger samples, for example, the expected dynamical evolution of multi-planet systems due to mutual perturbations (Teixeira & Ballard, 2023).666The orbital period distributions of the young and old sub-samples presented here do not appear significantly different (K-S test ).
| Name | Disposition | Period | Radius | Age |
| days | Gyr | |||
| KOI-247.01 | confirmed | 13.82 | 1.91 +0.10/-0.11 | 1.030.35 |
| KOI-251.01 | confirmed | 4.16 | 2.65 +0.10/-0.10 | 1.090.20 |
| KOI-251.02 | confirmed | 5.77 | 0.85 +0.05/-0.05 | 1.090.20 |
| KOI-253.01 | confirmed | 6.38 | 3.20 +0.13/-0.13 | 6.000.97 |
| KOI-253.02 | confirmed | 20.62 | 1.85 +0.10/-0.11 | 6.000.97 |
| KOI-254.01 | confirmed | 2.46 | 11.54 +0.35/-0.35 | 1.140.24 |
| KOI-314.01 | confirmed | 13.78 | 1.50 +0.05/-0.05 | 2.210.67 |
| KOI-314.02 | confirmed | 23.09 | 1.33 +0.05/-0.05 | 2.210.67 |
| KOI-314.03 | confirmed | 10.31 | 0.64 +0.03/-0.03 | 2.210.67 |
| KOI-430.01 | confirmed | 12.38 | 2.69 +0.10/-0.10 | — |
@@ Acknowledgements
Appendix A Biases and Selection Effects
We evaluated potential biases and selection effects that could cause the reconstruction of the planet radius distribution to mimic the observed evolution.
Choice of age divisor: We investigated the sensitivity of the difference in radius distribution between young and older systems to the choice of age to divide the sub-samples. The choice of median (3.86 Gyr, horizontal line in the top panel of Fig. 10) is a statistical convenience and gyrochronologic ages have significant uncertainties which can “blur" age dependence. To assess this, we generated alternative radius distributions, dividing the sample a the 40 and 60 percentile values of the age distribution (3.65 and 4.59 Gyr). These bound the grey shaded region in the top panel of Fig. 10. In the bottom panel, the radius distributions for the younger and older sub-samples are shown as the red and blue curves, respectively, with the solid curves those constructed using the median age and the shaded regions bounded by the results from the 40 and 60 percentile values. All younger distributions exhibit an elevated number of sub-Neptunes to super-Earths whereas for older distributions the reverse is true. Thus the evidence for radius evolution is robust to the exact choice of sample divisor.
Planet detection bias: We also investigated whether the observed radius evolution could be an artefact of differences in planet detection efficiency between young and old stars777The main sequence lifetime of stars increases with decreasing stellar mass but this is a consideration only for host stars more massive than the Sun.. Rotation-driven magnetic activity (spots and flaring) on cool stars introduces noise into lightcurves. As stars age and spin down, the noise decreases (e.g., Davenport, 2016), meaning that, all else being equal, smaller and/or more distant planets can be detected around older stars. To evaluate the magnitude of this systematic, we divided the Kepler late K/early M dwarf sample with rotation periods in Santos et al. (2019) into 773 (older) stars more slowly rotating than the rotation sequence of the 4 Gyr-old cluster M67 (Dungee et al., 2022), and 1936 (younger) stars that are more rapidly rotating than that sequence. The age of M67 is within error of our median sample age (3.86 Gyr). We calculated overall detection efficiency among the two samples for each planet in our KOI list and plot the ratio vs. planet radius in Fig. 11. All planets in the radius range of interest are less detectable by about 10% around the more rapidly rotating (“younger") stars than around the more slowly rotating (“older") stars and the difference is largest among the smallest planets. However, the relative decline in detection efficiency between super-Earths and sub-Neptunes is 10% and cannot explain the factor of 2 differences between the younger and older radius distributions (Fig. 5). Because we cannot adequately filter the overall sample for binary stars, which tend to be more rapidly rotating than single stars and also dilute any transit signals, we do not explicitly correct for this difference in detection efficiency.
Rotation period selection bias: This observed change in planet radii could possibly arise from a difference in the properties of younger and older stars brought about by selection for detectable rotation periods. Cooler dwarf stars tend to rotate more slowly at a given age (e.g., Curtis et al., 2019) and their rotation is more difficult to detect due to the longer time baseline needed to capture variability as well as the lower amplitude of variability from decreased magnetic activity. Thus, cooler stars are more likely to be included in the younger sub-sample, an effect which is apparent in the distribution of Kepler target stars (not just KOIs) with rotation-based ages younger or older than 3.86 Gyr (Fig. 12). This bias with should also bias the distributions with stellar radius, and potentially detection of transiting planets. However, we find that this does not produce any significant difference in planet detectability. The distribution of predicted transit signal-to-noise ratio (SNR) of the two subsets of target stars are nearly indistinguishable (Fig. 13). The median values are 0.35 and 0.36 and a two-sided Kolmogorov-Smirnov (K-S) returns a probability of 0.13 that the two samples, excluding the tails with relative SNR 1.3, are drawn from the same distribution. Here, we calculated the relative SNR for a given and as the inverse of the stellar radius times the square root of the brightness times the transit duration, which scales as , where Gaia magnitude was used as a proxy for Kepler magnitude . Cooler dwarf stars are smaller but also intrinsically fainter, and these two opposing trends apparently cancel out in this field.
Planet population variation: A dependence of the intrinsic planet population on stellar mass, combined with a difference in the mass distribution between younger and older sub-samples, could mimic evolution. The occurrence ratio of sub-Neptunes to super-Earths to sub-Neptunes modestly decreases with stellar mass between 0.5 and 1.4M⊙ (about 0.04 per 0.1 dex in mass, top panels of Fig. 12 in Petigura et al., 2022), but seems to markedly increase with increasing stellar mass in the M dwarf range (Hardegree-Ullman et al., 2019; Ment & Charbonneau, 2023). Thus, if there was a bias towards more higher-mass stars in the older sample due to a detection bias for shorter rotation periods (see above), the older sample should contain more sub-Neptunes, opposite to what is observed. Moreover, there is no apparent trend between host star luminosity (here , a proxy for mass) and age in our sample (Fig. 14). A two-sided K-S test returns a probability of 0.29 that the luminosities of younger and older stars are drawn from the same distribution.